Mr.Ahsan has an income of Rs.50,000 and he purchases 200 units of Y-GOOD. His
income increases by 30% and now he can purchase 220 units of Y-GOOD. Calculate
elasticity.
1
Expert's answer
2020-05-26T11:29:10-0400

"E=\frac{\Delta Demand }{\Delta Income}"

The change in Income - 30%,

the change in demand "\frac{220-200}{200}=10" %.

Thus, "E=\frac{10}{30}=1/3."